OpenStudy (anonymous):

A right circular cone is inscribed in a hemisphere. What is the ratio of the volume of the cone to the volume of the hemisphere? Volume of cone = 1/3πr^2h Volume of hemisphere = 2/3πr^3

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\pi r^2h \div \frac{ 2 }{ 3 } \pi r^3=\frac{ 1 }{\cancel{ 3 }}\cancel {\pi r^2}h \times \frac{ \cancel{3}}{ 2 \cancel{(\pi r^2)}r }=\frac{ h }{ 2r }\] @prettyboy20202

OpenStudy (anonymous):

could you explain this verbally so i can better understand?